OpenStudy (anonymous):

how do you find the solution set of x^2+25=0?? i know the answer is {5i,-5i} but not sure why

OpenStudy (anonymous):

subtract \(25\) from both sides, then take the square root

OpenStudy (anonymous):

i did but i dont know why the imaginary numbers are included in the answer

OpenStudy (anonymous):

you get \(x^2=-25\) so \(x=\pm\sqrt{-25}\)

OpenStudy (anonymous):

all you are doing is rewriting \(\sqrt{-25}+\sqrt{25\times (-1)}=\sqrt{25}\sqrt{-1}=5\sqrt{-1}\)

OpenStudy (anonymous):

and then rewriting \(\sqrt{-1}\) as \(i\) which is why it is \(5i\)

OpenStudy (anonymous):

i understand it now! thank you so much! :D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

wait... theres 5i but how do get -5i??

OpenStudy (anonymous):

because of the + or -?

OpenStudy (anonymous):

oh because it is exactly the same is if you had solve \(x^2-25=0\) you get \(x^2=25\) and so \(x=5\) or \(x=-5\) always two solutions

OpenStudy (anonymous):

in this case you get \(x^2=-25\) so \(x=5i\) or \(x=-5i\) as above

OpenStudy (anonymous):

like -5x5=25 and 5x5 =25?

OpenStudy (anonymous):

*-5x-5=25

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

in this case \((-5i)(-5i)=25i^2=-25\)

OpenStudy (anonymous):

thanks again :)